OpenStudy (cyrusleaf1101):

Write the equation of the line that passes through (–1, 5) and has a slope of 3 in point-slope form? x + 1 = 3(y – 5) x – 5 = 3(y + 1) y – 5 = 3(x + 1) y + 1 = 3(x – 5)

OpenStudy (visarika23):

Eq of line in point slope form is (y-y1)=m(x-x1) Where (x1,y1) is the point and m is slope Here x1=-1 y1=5 and m=3 Substitute these values and get the eq

OpenStudy (cyrusleaf1101):

so it would be a? x+1=3(y-5) i am terrible at math so idk :/

OpenStudy (visarika23):

No it wil be (y-5)=3(x+1)

OpenStudy (cyrusleaf1101):

ty :) you are a great helper :)

OpenStudy (visarika23):

Yr wlcm☺